# HA-VLN: A Benchmark for Human-Aware Navigation in Discrete–Continuous Environments with Dynamic Multi-Human Interactions, Real-World Validation, and an Open Leaderboard

### v1

@@ 4 HA-VLN Agents
This section introduces our Human-Aware Room-to-Room (HA-R2R) dataset and two navigation agents: HA-VLN-VL and HA-VLN-CMA. We aim to endow agents with enhanced understanding of social cues—respecting personal space, avoiding collisions, and reacting to bystanders. As illustrated in Figures A8 and A9, dynamic human behaviors impose additional complexity and variability, necessitating robust perception and policy strategies.
@@
HA-R2R Dataset. Although the Room-to-Room in Continuous Environment (R2R-CE) dataset [28] supports continuous navigation, it lacks explicit focus on human interactions. We therefore develop HA-R2R, which extends R2R-CE with 16,844 carefully curated instructions highlighting social nuances (e.g., conversations, corridor crossings, near-collision events). Table A4 showcases representative directives, while Figure A6 reveals the expanded vocabulary centered around human-centric language.
@@
To generate these enriched instructions, we design targeted prompts for large language models (Appendix C.2), capturing a diverse range of real-world social scenarios. This augmentation marks a significant leap from static paths to ones requiring agents to interpret, for example, “avoid the couple chatting near the bar” or “yield to someone crossing the hallway.” Our comparative analysis between baseline and augmented instructions (Appendix C.3) highlights both the extensive annotation workload and the potential of HA-R2R to advance human-aware navigation research.
HA-VLN-VL Agent. HA-VLN-VL agent focuses on visual-language alignment when encountering socially enriched instructions. Adapted from Recurrent VLN-BERT [20], it discards actor-critic methods (e.g., A2C [27]) in favor of a simpler imitation learning approach, thereby emphasizing how stronger multimodal grounding alone can elevate performance in complex settings.
At each timestep $t$, the agent updates its hidden state $s_{t}$ and predicts an action distribution $p_{t}^{a}$:
$s_{t},\;p_{t}^{a}\;=\;\text{HA-VLN-VL}(s_{t-1},\,X,\,V_{t}),$
@@
where $X$ is the tokenized instruction (often referencing multiple humans or group interactions), and $V_{t}$ encodes the fused RGB-depth view. A multi-layer Transformer processes these inputs, using a specialized state token to attend to language and visual tokens. Final action probabilities result from an average of attention weights:
$p_{t}^{a}\;=\;\overline{\text{AveragePool}}_{s,v}^{l}.$
By fine-tuning from Prevalent [20] on HA-R2R, HA-VLN-VL concentrates on the semantic interplay between social directives and the observed scene. Further architectural details appear in Appendix C.6.
@@
HA-VLN-CMA Agent. While HA-VLN-VL excels at interpreting nuanced instructions, HA-VLN-CMA prioritizes collision avoidance & real-time adaptation in human-filled environments. Drawing on cross-modal attention (CMA) [28], it fuses textual embeddings $l=\text{BERT}(I)$ and visual features $v_{t}=\text{ResNet}(o_{t})$ each timestep. A multi-head attention mechanism produces a combined representation $f_{t}$, which an MLP maps to:
$P(a_{t}\mid f_{t})\;=\;\text{Softmax}\!\bigl{(}\text{MLP}_{\mathrm{action}}(f_{t})\bigr{)}.$
Figure A7(b) outlines the pipeline, while Appendix C.7 discusses implementation details. Human-populated settings entail partial observability and unpredictable motion. To address these challenges, we leverage Environmental Dropout (Envdrop) [46] & Dataset Aggregation (DAgger) [44]. Envdrop randomly masks features in the visual-language streams, simulating occlusions by crowds or obstructing objects. DAgger maintains a replay buffer of agent states under both expert and learned policies, enabling the agent to iteratively correct prior errors. This additional training effort improves trajectory re-planning when confronted with obstacles or unexpected human movements.
Algorithm 1 HA-VLN-CMA Training Procedure
0:  Expert policy $\pi^{*}$, parameters $\theta$, replay buffer $B$
0:  Trained CMA model parameters $\theta$
1:  Initialize parameters $\theta$ (e.g., via imitation learning)
2:  for iteration $=1$ to $N$ do
3:     Execute policy $\pi_{\theta}$ to collect rollouts
4:     With probability $\beta$, follow expert $\pi^{*}$
5:     Store $(s_{t},a_{t},r_{t})$ in $B$
6:     Apply Envdrop to $(v_{t},l)$
7:     $\theta\leftarrow\theta-\eta\nabla_{\theta}\,\mathcal{L}_{\mathrm{DAgger}}(B,\theta)$
8:  end for
Analysis & Insights. Figures A8 & A9 demonstrate how our agents respond to bystanders, underscoring that collision likelihood and route deviation both significantly increase when humans crowd key passageways. HA-VLN-CMA re-plans aggressively if a person suddenly blocks its immediate path, whereas HA-VLN-VL draws upon textual grounding to maintain a polite distance from social gatherings. This distinction showcases our dual innovation: a socially enriched dataset (HA-R2R) and specialized models that jointly facilitate robust human-aware navigation. Table A4 & Figure A6 further illustrate the magnitude and complexity of our annotations, emphasizing the extensive workload in data preparation, model adaptation, and thorough validation. Finally, in Section 6, we present comprehensive evaluations on continuous & discrete tasks derived from HA-R2R, illustrating the complementary strengths of HA-VLN-VL (strong language-vision fidelity) and HA-VLN-CMA (robust collision avoidance).
Figure 5: Human-aware navigation with multiple bystanders. Top: The robot receives instructions emphasizing caution as one person moves from the lounge to the plant while another prepares to exit. Middle: A third-person view highlights the robot’s trajectory and path planning among dynamic bystanders. Bottom: The onboard camera shows the robot’s egocentric observations.
5 Real-World Validation & Leaderboard
Real-World Validation & Setup. We deploy our trained agents on a Unitree Go2-EDU quadruped, equipped with an Intel Realsense D435i RGB-D camera, MID360 3D LiDAR, and IMU for onboard perception and control. As Figure 5 illustrates, experiments are conducted in four indoor spaces (office, living room, hallway, lobby), each populated by 2–4 free-moving volunteers. The agent navigates safely in moderately congested conditions but faces challenges in tight corridors or sudden crowd convergence, highlighting the need for robust re-planning under partial observability.
HA-R2R Test Dataset & Leaderboard. Building on R2R-CE, we present HA-R2R, featuring 16,844 instructions across 90 building scans with 910 annotated human models (see Sections 34). The test partition—3,408 instructions in 18 withheld buildings—intentionally emphasizes multi-human routes. While retaining path continuity from R2R-CE, we introduce refined goals to stress social awareness. The dataset is split into training (10,819), seen validation (778), unseen validation (1,839), and testing (3,408). We also host leaderboards for HA-R2R-DE (discrete) and HA-R2R-CE (continuous), focusing on both collision-related (TCR, CR) and navigation (NE, SR) metrics. Submissions may include agent code or trajectories, providing reproducible, server-side evaluations and setting a new benchmark for human-centric, dynamic VLN research.
@@ 5 Real-World Validation & Leaderboard
Table 1: HA-VLN-CE Results Across Validation (Seen/Unseen) and Test Splits. “HA-VLN-CMA” denotes the full version of HA-VLN-CMA (+DA +EV). Metrics include NE (Navigation Error, meters), TCR (Total Collision Rate), CR (Collision Rate per step), and SR (Success Rate), with lower NE/TCR/CR and higher SR indicating better performance.
@@
BEVBert [1] 5.53 3.64 0.46 0.27 6.11 4.29 0.47 0.19 5.51 4.71 0.55 0.21 6.10 5.72 0.56 0.15 6.33 4.25 0.58 0.18 6.54 4.39 0.54 0.14
ETPNav [2] 5.17 4.07 0.43 0.24 7.72 6.31 0.61 0.12 5.43 6.94 0.58 0.17 7.40 7.94 0.71 0.08 5.94 5.96 0.58 0.16 7.59 5.64 0.73 0.09
Table 2: Performance in Unseen Environments (HA-VLN-CE) with Relative Gaps. We highlight $\dagger$” for zero-shot models (trained on VLN-CE). Columns for NE/TCR/CR/SR are available in Table 1, so here we focus on relative improvements (%). Entries show changes vs. the VLN-CE baseline.
Agent Retrained Gap (%) Zero-Shot Gap (%)
[…]
HA-VLN-CMA-Base -9.8 -69.6
HA-VLN-CMA-DA -13.8 -66.7
HA-VLN-CMA -10.0 -68.8 +10.2 -71.9
HA-VLN-VL -6.8 -68.2 +24.6 -86.4
BEVBert [1] +20.6 -64.4 +33.5 -74.6
ETPNav [2] +15.3 -70.2 +57.1 -86.0
Table 3: Retrained Performance on HA-VLN-DE (Unseen Environments). NE: Navigation Error (meters, lower is better), SR: Success Rate (higher is better), TCR: Total Collision Rate (lower is better), CR: Collision Rate per Step (lower is better).
@@
Speaker-Follower [12] 6.62 0.35 7.44 0.32 0.72 0.21
Rec (PREVALENT) [20] 3.93 0.63 6.12 0.29 0.81 0.33
Rec (OSCAR) [20] 4.29 0.59 6.12 0.28 0.78 0.34
Airbert [16] 4.01 0.62 5.54 0.30 0.83 0.36
Table 4: Impact of Human Presence (hp) and Interaction Enrichment (enrich). We evaluate both continuous (CE) and discrete (DE) settings, reporting NE: Navigation Error, TCR: Total Collision Rate, CR: Collision Rate per step, SR: Success Rate.
@@
DE Rec (PREVALENT) 7.52 ($\uparrow$2.9%) 0.27 ($\downarrow$12.9%) 0.64 ($\downarrow$19.0%) 0.27 ($\uparrow$22.7%)
@@
DE Rec (PREVALENT) 7.33 ($\uparrow$0.3%) 0.19 ($\downarrow$38.7%) 0.42 ($\downarrow$46.8%) 0.26 ($\uparrow$18.2%)
Table 5: Cross-Evaluation of BEVBert (CE) vs. Rec (PREVALENT) (DE). Each model is trained/validated under two different instruction sets (HA-R2R vs. R2R/R2R-CE). NE: Navigation Error (m, lower is better); SR: Success Rate (higher is better). The gray cells () indicate performance changes when models are trained on R2R/R2R-CE instructions but validated on HA-R2R.
Env Training Validation Val (Unseen)
[…]
CE HA-VLN-CE HA-R2R HA-VLN-CE HA-R2R 5.51 0.21
R2R-CE HA-R2R 6.23 ($\uparrow$13.1%) 0.15 ($\downarrow$28.6%)
DE HA-VLN-DE HA-R2R HA-VLN-DE HA-R2R 5.01 0.39
R2R HA-R2R 6.11 ($\uparrow$22.0%) 0.24 ($\downarrow$38.5%)
Table 6: Impact of Step Size on Navigation. Increasing the step size transitions from finer-grained (CE-like) control to more discrete steps. We show results for both BEVBert [1] and ETPNav [2] on seen/unseen validations. NE: Navigation Error (m), TCR: Total Collision Rate, CR: Collision Rate per step, SR: Success Rate. The default step size for CE is 0.25 m.
@@
Table 7: Ablation on RGB/Depth Inputs. We compare BEVBert [1] and ETPNav [2] on seen/unseen validations. denotes the sensor is enabled, while is disabled. Gray cells highlight relative performance changes (in %) upon removing or adding a modality. NE: Navigation Error (m), TCR: Total Collision Rate, CR: Collision Rate per step, SR: Success Rate. Best viewed in color.
@@
BEVBert[1] 6.23 ($\uparrow$12.6%) 4.55 ($\uparrow$25.0%) 0.49 ($\uparrow$6.5%) 0.19 ($\uparrow$29.6%) 5.79 ($\uparrow$5.1%) 4.97 ($\uparrow$5.5%) 0.53 ($\uparrow$3.6%) 0.15 ($\downarrow$28.6%)
@@
ETPNav[2] 6.14 ($\uparrow$18.8%) 6.07 ($\uparrow$49.1%) 0.56 ($\uparrow$30.2%) 0.17 ($\uparrow$29.2%) 6.38 ($\uparrow$17.5%) 7.44 ($\uparrow$7.2%) 0.65 ($\uparrow$12.1%) 0.13 ($\downarrow$23.5%)
@@
Table 8: Step Size Combination Performance Across Scenarios. NE: Navigation Error (m), TCR: Total Collision Rate, CR: Collision Rate per step, SR: Success Rate.
[…]
1.00 6.85 26.97 0.94 0.004
2.25 8.79 112.78 0.97 0.000
Table 9: Navigation Success Rate (NSR, %) in four environments (Living Room, Office, Hallway, and Lobby) with or without human presence (w/o vs. w/), averaged over 10 trials. Bold numbers indicate the highest performance in each column.
Method Living Rm Office Hallway Lobby Overall
w/o w/ w/o w/ w/o w/ w/o w/ w/o w/
HA-VLN-CMA-Base 0.2 0.1 0.2 0.1 0.2 0.1 0.2 0.1 0.2 0.1
HA-VLN-VL 0.4 0.1 0.4 0.1 0.5 0.2 0.4 0.1 0.4 0.1
6 Experiments
We evaluate our methods in two primary settings: (i) human-populated simulations within HA-VLN-CE (continuous) and HA-VLN-DE (discrete), and (ii) real-world robot deployments. All evaluations use four metrics—Navigation Error (NE), Success Rate (SR), Total Collision Rate (TCR), and Collision Rate (CR)—to measure navigation accuracy and social compliance. Extended formulas appear in Appendix D.1.
6.1 HA-VLN-CE: Continuous Navigation
Task Definition. HA-VLN-CE extends VLN-CE [28] by populating photorealistic 3D environments with multiple, independently moving humans. Agents navigate continuously (forward, turn) based on natural-language instructions, aiming to reduce NE while limiting collisions (TCR, CR). This setup mirrors real-world conditions where bystanders can alter their paths unexpectedly, demanding reactive policies and sophisticated sensory integration.
Baselines. We systematically benchmark two notable continuous-navigation models, BEVBert [1] and ETPNav [2], together with our HA-VLN-CMA and HA-VLN-VL agents. Each approach is evaluated under two configurations: Retrained, which leverages HA-R2R’s multi-human data, and Zero-shot, which relies solely on R2R-CE training. Contrasting these configurations further underscores the importance of explicit, human-centric instruction for truly robust social navigation.
Results. Table 1 shows pronounced gains when models incorporate HA-R2R data. For instance, BEVBert’s TCR drops from 4.29 to 3.64 in seen environments and from 5.72 to 4.71 in unseen ones. In stark contrast, Zero-shot models frequently collapse in high-density scenarios, with TCR surpassing 30 if they have no exposure to human-centric interactions. Table 2 further demonstrates that omitting multi-human data can diminish SR by up to 70% in unseen splits, underscoring the critical role of socially aware training for both collision avoidance and trajectory completion. These improvements suggest that explicit references to bystanders and dynamic crowd behavior substantially enhance real-world navigational readiness.
Ablation Studies. (1) Discrete vs. Continuous. Table 3 finds that discrete agents (HA-VLN-DE) can achieve moderate SR yet suffer high collisions in crowded scenes, implying simpler viewpoint-based control is inadequate for dynamic social contexts. (2) Human Presence. Table 4 shows how higher bystander density escalates TCR, underscoring the need for more robust collision avoidance in large crowds. (3) Instruction Alignment. Table 5 reveals that replacing HA-R2R with standard R2R-CE text raises NE by 13–22% and cuts SR by up to 38%, confirming the value of explicit human-motion references. (4) Step Size. Tables 6 and 8 show finer increments refine crowd responses but may inflate TCR via excess micro-adjustments. (5) Sensor Modalities. Table 7 confirms that adding depth consistently lowers collisions and raises SR, reflecting the importance of 3D cues for navigating around moving bystanders.
Insights. Together, these findings confirm that continuous, human-aware navigation benefits significantly from training on multi-human data, realistic instructions referencing bystander behavior, and richer sensor inputs. Failure to address any of these factors can lead to more collisions and lower success, highlighting the multi-faceted challenges of real-world social navigation.
6.2 HA-VLN-DE: Discrete Navigation.
Task Definition. HA-VLN-DE extends panoramic VLN [3] by placing multiple moving individuals in a viewpoint-based environment. Although discrete control restricts possible movements to predefined locations, it does not eliminate collisions, as bystanders may occupy a chosen viewpoint or block key transitions in narrow corridors.
Results. Table 3 compares top discrete agents on multi-human tasks. While Airbert achieves a moderate SR=0.36, it can incur a CR up to 0.83, illustrating persistent collision risks. Approaches that overlook human dynamics or personal space often fail when multiple bystanders converge, particularly in tight junctions or doorways. Consequently, viewpoint-based navigation alone is insufficient for robust social navigation—adaptive collision-avoidance strategies remain essential even in discrete settings.
6.3 Real-World Experiments.
Setup. We evaluate our agents in four everyday indoor environments—office, living room, hallway, and lobby—under two conditions: (i) no bystanders and (ii) 2–4 free-moving volunteers. This setup simulates realistic indoor traffic patterns and partial observability.
Observations. As illustrated in Figure 5, the robot frequently pauses or yields to avoid oncoming pedestrians. In the absence of bystanders, it navigates smoothly (Figure A11), but collisions arise in cramped corridors or when crowds converge suddenly (Figure A12). We observe similar patterns in living-room environments (Figures A13A14) and hallways (Figure A15).
Table 9 shows the average NSR (Navigation Success Rate) across 10 trials in each environment. While human presence invariably lowers NSR, HA-VLN-VL consistently outperforms HA-VLN-CMA-Base, demonstrating stronger adaptability to dynamic motion. Still, partial observability and abrupt group formations remain challenging, especially in narrow passages or at congested junctions. Additional metrics—such as collision frequencies and path divergence—appear in Appendix D.4, where we further detail performance under varying crowd densities.
Overall, these experiments confirm that simulation-trained, multi-human navigation policies can indeed transfer to physical robots. However, further refinement in collision forecasting and reactive control is needed to handle unpredictable human behavior in tight indoor settings.

### v3

@@ 4 HA-VLN Agents
To ground the unified HA-VLN task in our HA-VLN simulator, we introduce the Human-Aware Room-to-Room (HA-R2R) dataset and two baseline agents, HA-VLN-VL and HA-VLN-CMA. These agents are designed as reference implementations rather than final solutions, offering a starting point for developing more advanced models. They emphasize essential social capabilities including maintaining personal space, avoiding collisions, and adapting to bystanders, under the dynamic conditions of HA-VLN 2.0. As shown in Figs. A8 and 5, human behaviors add substantial complexity, motivating the dataset design and agent baselines described in the following paragraphs.
@@
HA-R2R Dataset. The Room-to-Room in Continuous Environment (R2R-CE) dataset Krantz et al. (2020) supports continuous navigation but lacks explicit modeling of human interactions. We therefore extend it into HA-R2R, which contributes 16,844 curated instructions emphasizing social nuance, covering conversations, corridor crossings, and near-collision events. Table A3 presents representative directives, while Fig. A6 visualizes the expanded human-centric vocabulary.
Figure 4: HA-R2R Dataset Analysis. (a) Distribution of instruction length by human group size (none to >>3). (b) Comparison of instruction lengths between HA-R2R and R2R-CE.
@@
We generate these enriched instructions via targeted LLM prompts (Appendix C.2), capturing diverse social scenarios. This augmentation shifts navigation from static paths to socially contingent routes, e.g., “avoid the couple chatting near the bar.” Comparative analyses (Appendix C.3) highlight both the annotation workload and HA-R2R’s potential for human-aware navigation.
HA-VLN-VL Agent. The HA-VLN-VL focuses on visual–language alignment. Adapted from Recurrent VLN-BERT Hong et al. (2021), it replaces actor–critic methods (e.g., A2C Konda & Tsitsiklis (1999)) with a streamlined imitation learning objective, isolating the contribution of multimodal grounding. At timestep $t$, the agent updates its hidden state $s_{t}$ and predicts an action distribution:
$s_{t},\;p_{t}^{a}\;=\;\text{HA-VLN-VL}(s_{t-1},\,X,\,V_{t}),$ (3)
@@
where $X$ is the tokenized instruction (often referencing multiple humans) and $V_{t}$ encodes the fused RGB–depth view. A Transformer with a specialized state token attends jointly to visual and linguistic tokens, and final probabilities are derived via pooled attention:
$p_{t}^{a}\;=\;\overline{\text{AveragePool}}_{s,v}^{l}.$ (4)
Fine-tuned from Prevalent Hong et al. (2021) on HA-R2R, HA-VLN-VL demonstrates how stronger grounding alone benefits navigation under socially complex conditions (Appendix C.6).
@@
HA-VLN-CMA Agent. HA-VLN-CMA emphasizes collision avoidance and real-time adaptation. Built on cross-modal attention (CMA) Krantz et al. (2020), it fuses textual embeddings $l=\text{BERT}(I)$ with visual features $v_{t}=\text{ResNet}(o_{t})$. Multi-head attention produces a joint representation $f_{t}$, which an MLP maps to action probabilities:
$P(a_{t}\mid f_{t})\;=\;\text{Softmax}\!\bigl(\text{MLP}_{\mathrm{action}}(f_{t})\bigr).$ (5)
Fig. A7(b) outlines the architecture (details in Appendix C.7). To address partial observability and unpredictable motion, we adopt Environmental Dropout (Envdrop) Tan et al. (2019) to simulate occlusions and Dataset Aggregation (DAgger) Ross et al. (2011) for iterative error correction. These strategies enhance re-planning when agents face obstacles or unexpected behaviors. Figs. A8 and 5 illustrate agent responses to bystanders, showing that collision risk and route deviation increase sharply in crowded passages. HA-VLN-CMA re-plans aggressively when blocked, whereas HA-VLN-VL leverages textual grounding to maintain appropriate distances. This contrast highlights our dual contributions: a socially enriched dataset (HA-R2R) and two baseline agents serving as extensible reference points. These baselines are not final solutions but starting points for the community to build, refine, and extend toward more advanced human-aware navigation models. Sec. 5 evaluates both agents on HA-VLN 2.0, demonstrating complementary strengths.
5 Experiments
Evaluation Metrics. We evaluate performance on the HA-VLN 2.0 benchmark using two suites of metrics. (1) Social compliance. To assess social awareness, we use Total Collision Rate (TCR) and Collision Rate (CR). TCR measures the overall frequency of collisions, while CR reflects the proportion of socially inappropriate interactions. (2) Navigation accuracy. We report Navigation Error (NE) and Success Rate (SR). A trajectory is deemed successful under SR not only when the agent stops sufficiently close to the goal Anderson et al. (2018), but also when it demonstrates effective obstacle avoidance. Formal definitions of these metrics are provided in Appendix D.1.
@@ 5 Experiments
Table 1: HA-VLN-CE Results Across Validation (Seen/Unseen) and Test Splits. “HA-VLN-CMA” denotes the full version of HA-VLN-CMA (+DA +EV). Metrics include NE (Navigation Error, meters), TCR (Total Collision Rate), CR (Collision Rate per step), and SR (Success Rate), with lower NE/TCR/CR and higher SR indicating better performance. All agents receive panoramic RGBD observations at each location.
@@
BEVBert An et al. (2023) 5.53 3.64 0.46 0.27 6.11 4.29 0.47 0.19 5.51 4.71 0.55 0.21 6.10 5.72 0.56 0.15 6.33 4.25 0.58 0.18 6.54 4.39 0.54 0.14
ETPNav An et al. (2024) 5.17 4.07 0.43 0.24 7.72 6.31 0.61 0.12 5.43 6.94 0.58 0.17 7.40 7.94 0.71 0.08 5.94 5.96 0.58 0.16 7.59 5.64 0.73 0.09
Figure 5: Agent Trajectory Examples (HA-VLN-CMA). The top row demonstrates a failed navigation scenario where the agent fails to avoid an oncoming human, ultimately resulting in a collision. In contrast, the bottom row showcases a successful navigation: the agent proactively adjusts its trajectory to the left, effectively avoiding human interference and completing the task without collision.
We evaluate agents in two settings: (1) We assess the performance of HA-VLN 2.0 agents alongside several top agents on the HA-VLN 2.0 benchmark, utilizing both HA-VLN-CE (continuous) and HA-VLN-DE (discrete) (Sec. 5.1). We conduct extensive analysis and ablation studies examining key factors including continuous versus discrete settings, cross-domain generalization capabilities, human presence and interaction enrichment, step size variations, and sensor modality configurations. These analyses investigate their respective impacts on human-aware navigation performance and reveal complementary knowledge between the DE and CE approaches. (2) We deploy and evaluate HA-VLN 2.0 agents in real-world robotic scenarios across diverse layouts (office spaces, living rooms, hallways, and lobbies) with free-moving human volunteers (Sec. 5.2, Appendix D.5).
5.1 Benchmarking Agents on HA-VLN 2.0
HA-VLN-CE. We systematically benchmark two notable continuous navigation agents, BEVBert An et al. (2023) and ETPNav An et al. (2024), together with our HA-VLN-CMA and HA-VLN-VL agents in Table 1. Each approach is trained/evaluated under two configurations: Retrained, where agents are trained/evaluated solely on HA-VLN-CE benchmark (HA-VLN-CE simulator + HA-R2R instruction dataset), and Zero-shot, where agents are trained solely on VLN-CE benchmark (VLN-CE simulator + R2R-CE) and evaluated on our benchmark. Table 1 shows pronounced gains when models incorporate HA-VLN-CE benchmark. For instance, BEVBert’s SR increases from 0.19 to 0.27 in seen split and from 0.15 to 0.21 in unseen. In stark contrast, Table 3 shows that BEVBert trained on our benchmark performs comparably to the VLN-CE-trained one on VLN-CE benchmark (SR: 0.35 vs. 0.37). This bidirectional evaluation suggests that explicit references to dynamic crowd behavior enhance real-world navigational readiness and confirm the robustness of HA-VLN-CE. Figure 5 presents navigation visualizations of HA-VLN-CMA agent on the HA-VLN-CE benchmark, including one successful and one failed example. These examples demonstrate that dynamic human activities indeed increase the difficulty of navigation, while also making the scenarios more realistic and reflective of real-world challenges.
HA-VLN-DE. Table 2 compares top discrete agents on both VLN and HA-VLN-DE benchmarks, showing discrete agents can achieve moderate SR yet suffer high collisions in crowded scenes.
Table 2: DE performance of agents trained on VLN vs. HA-VLN-DE (Unseen). All agents use panoramic RGB observations.
@@
Speaker-Follower 6.62 0.35 7.44 0.32 0.72 0.21
Rec (PREVALENT) 3.93 0.63 6.12 0.29 0.81 0.33
Rec (OSCAR) 4.29 0.59 6.12 0.28 0.78 0.34
Airbert 4.01 0.62 5.54 0.30 0.83 0.36
NavCoT 6.26 0.40 6.83 0.36 0.85 0.23
For example, while Airbert Guhur et al. (2021) achieves a moderate SR at 0.36, it can incur a CR of up to 0.83, illustrating persistent collision risks. The results showcase adaptive collision-avoidance strategies also remain essential in discrete settings. Approaches that overlook human dynamics often fail when multiple bystanders converge (Sec. 3), particularly in tight junctions or doorways.
Table 3: Cross Domain Evaluation of BEVBert (CE) vs. Rec (PREVALENT) (DE). Each model is trained/validated under different simulators (HA-VLN-CE/HA-VLN-DE vs. VLN-CE/VLN-DE) and different instruction sets (HA-R2R vs. R2R-CE/R2R). The blue cells () indicate performance changes when models are trained on R2R/R2R-CE instructions but validated on HA-R2R.
Env Training Validation Val (Unseen)
[…]
CE VLN-CE R2R-CE VLN-CE R2R-CE 4.57 0.37
HA-VLN-CE HA-R2R R2R-CE 5.11 0.35
HA-VLN-CE HA-R2R R2R-CE 4.35 0.27
R2R-CE R2R-CE 4.13 0.29
HA-R2R HA-R2R 5.51 0.21
R2R-CE HA-VLN-CE HA-R2R 6.23 ($\uparrow$13.1%) 0.15 ($\downarrow$28.6%)
Env Training Validation Val (Unseen)
[…]
DE VLN-DE R2R VLN-DE R2R 3.93 0.48
HA-VLN-DE R2R R2R 4.62 0.45
HA-VLN-DE HA-R2R HA-VLN-DE R2R 5.86 0.36
R2R R2R 5.21 0.33
HA-R2R HA-R2R 5.01 0.39
R2R HA-R2R 6.11 ($\uparrow$22.0%) 0.24 ($\downarrow$38.5%)
Table 4: Left: (a). Impact of Human Presence (hp) and Interaction Enrichment (enrich). We evaluate without hp (replace human with cylinders) and without enrich (skip interaction & movement enrichment in Sec. 3, Appendix B.5) on both CE and DE settings. Rec (PRE) denotes Rec (PREVALENT). Right: (b). Impact of Step Size on Navigation. Here the collision is detected only at endpoint of a step, thus increasing step size transitions from finer-grained control to more discrete (teleport-potential) steps (default step size for CE is 0.25m). We show results for both BEVBert An et al. (2023) and ETPNav An et al. (2024) on seen/unseen.
@@
DE Rec (PRE) 7.52 ($\uparrow$2.9%) 0.27 ($\downarrow$12.9%) 0.64 ($\downarrow$19.0%) 0.27 ($\uparrow$22.7%)
@@
DE Rec (PRE) 7.33 ($\uparrow$0.3%) 0.19 ($\downarrow$38.7%) 0.42 ($\downarrow$46.8%) 0.26 ($\uparrow$18.2%)
@@
Table 5: Ablation on RGB/Depth Inputs. We compare BEVBert An et al. (2023) and ETPNav An et al. (2024) on seen/unseen validations. denotes the sensor is enabled, while is disabled. Blue cells highlight performance changes (in %) upon removing/adding a modality. Best viewed in color.
@@
BEVBertAn et al. (2023) 6.23 ($\uparrow$12.6%) 4.55 ($\uparrow$25.0%) 0.49 ($\uparrow$6.5%) 0.19 ($\uparrow$29.6%) 5.79 ($\uparrow$5.1%) 4.97 ($\uparrow$5.5%) 0.53 ($\uparrow$3.6%) 0.15 ($\downarrow$28.6%)
@@
ETPNavAn et al. (2024) 6.14 ($\uparrow$18.8%) 6.07 ($\uparrow$49.1%) 0.56 ($\uparrow$30.2%) 0.17 ($\uparrow$29.2%) 6.38 ($\uparrow$17.5%) 7.44 ($\uparrow$7.2%) 0.65 ($\uparrow$12.1%) 0.13 ($\downarrow$23.5%)
@@
Analysis & Ablation Studies. (1) Cross-domain Generalization. Table 3 reveals that HA-R2R-trained agents achieve comparable SR to R2R-CE-trained agents (0.27 vs. 0.29) on R2R-CE validation set, while they outperform by +28.6% SR on the HA-R2R validation set, showcasing HA-R2R improves in-domain performance while maintaining cross-domain robustness. (2) Human Presence and Interaction Enrichment. Table 4 (a) shows in human presence ablations, replacing humans with cylinders drops TCR by around 36% and raises SR by around 10%, while removing human interaction enrichment drops TCR by up to 22% and raises SR by up to 25%, confirming humans are not merely treated as generic moving obstacles during navigation. (3) Step Size. Table 4 (b) indicates a degree of knowledge complementarity between DE and CE navigation when collisions are detected only at the endpoint of a step. Specifically, increasing the step size (from 0.1 m to 1.0 m), approximating DE-style navigation, can improve performance. We also conducted an additional experiment (Table A4) in which a 1.0 m step was treated as four 0.25 m sub-steps, and a 2.25 m step as nine 0.25 m sub-steps, with collisions checked after each sub-step. When evaluated on BEVBert in the val_unseen split, the agents failed to navigate effectively with both 1.0 m and 2.25 m step sizes, with SR dropping close to zero. These results highlight the need to account for the potentially “teleport-like” movement behaviors in DE when considering complementarity. (4) Sensor Modalities. Table 5 confirms that either adding depth or RGB consistently lowers collisions and raises SR, reflecting the importance of 3D cues for navigating around moving bystanders.
Figure 6: (a). Interactive interface we provide to explore 910 annotated human models and environments in HA-VLN 2.0 simulator from nine views. (b). Human-aware navigation with multiple bystanders. Left: Instruction provided to the robot. Right: A third-person view illustrates the robot’s trajectory among dynamic bystanders, and selected robot observations from onboard camera.
5.2 Leaderboard & Real-World Validation
HA-R2R Test Dataset & Leaderboard. Building on R2R-CE, we present HA-R2R, featuring 16,844 instructions across 90 building scans with 910 annotated human models (see Secs. 34). While retaining path continuity from R2R-CE, we introduce refined goals to emphasize social awareness. The test partition of HA-R2R contains 3,408 instructions across 18 withheld buildings and intentionally emphasizes multi-human routes. To assess performance on this challenging test split, we host leaderboards for HA-R2R-DE and HA-R2R-CE benchmarks, evaluating both collision-related metrics (TCR, CR) and navigation metrics (NE, SR). We prepare an interactive interface shown in Figure 6 (a), where participants can explore the simulator from nine different views to examine all the annotated human motions and the surrounding environments. This allows them to gain a deeper understanding of the challenging dynamic scenarios we provide. Submissions may include agent code or trajectories, providing reproducible, server-side evaluations and setting a new benchmark for human-centric, dynamic VLN research.
Real-World Validation & Setup. We deploy our trained agents on a Unitree Go2-EDU quadruped, equipped with Intel Realsense D435i RGB-D camera, MID360 3D LiDAR, and IMU for onboard perception and control. As Figure 6 (b) illustrates, experiments are conducted in four indoor spaces (office, living room, hallway, lobby), each populated by 2–4 free-moving volunteers. Implementation details and more visual examples are provided in Appendix D.5. The agent navigates safely in moderately congested conditions but faces challenges in tight corridors or sudden crowd convergence, highlighting the need for robust re-planning under partial observability.